OpenStudy (anonymous):

Eliminate the parameter. x = t^2 + 2, y = t^2 - 4

OpenStudy (anonymous):

Just do this... here 't' is the parameter and you have to eliminate it.... So, Why not subtract the equations ??? Or substitute value of any one out of 'x' and 'y' from one equation to the other... Hope this helps :-)

OpenStudy (anonymous):

so.... this? y = x - 6, x ≥ 1

OpenStudy (anonymous):

I think it's okay... pretty cool but I am not yet clear why you took x>or = 1... We can set anything we want for 'x' ...

OpenStudy (anonymous):

it is just in the answers idk why its like that lol

OpenStudy (anonymous):

'x' is the independent variable... so we can assign any value to 'x' ... depending on what 'x' is , y and t would have real or complex values... Anyways if that's the most APPROPRIATE out of the choices then it's fine...

OpenStudy (anonymous):

\[\begin{cases}x=t^2+2\\y=t^2-4\end{cases}\] Write both equations in terms of t²: \[\begin{cases}t^2=x-2\\t^2=y+4\end{cases}\] Now setting them equal to each other, you have \[x-2=y+4\\ y=x-6\]

OpenStudy (anonymous):

Quite the same thing ... braced with language :-)